OpenStudy (anonymous):

Can someone help me solve this natural log problem? ln10+ln5x^2=ln8

zepdrix (zepdrix):

\[\Large\bf\sf \ln(10)+\ln(5x^2)=\ln(8)\]We want to start by applying a rule of logs:\[\large\bf\sf \color{#A57F02}{\log(a)+\log(b)\quad=\quad \log(a\cdot b)}\]

OpenStudy (anonymous):

so ln10+ln5x^2=ln(10*5x^2) ??

zepdrix (zepdrix):

Ok good. Let's multiply the 10 and 5,\[\Large\bf\sf \ln(50x^2)=\ln(8)\]

zepdrix (zepdrix):

Apply the inverse of the log function to get rid of the logs on each side. Or just recognize that both sides are the `same log`, so their contents must be equal.\[\Large\bf\sf 50x^2=8\]

OpenStudy (anonymous):

ahhh now thats where I know how to solve, thank you for showing me the steps!! :)

zepdrix (zepdrix):

One thing to be careful of, Taking the square root of x^2 will give you `two solutions`, plus/minus. Since we're dealing with logs, you have to careful check and see if both solutions would work. (Meaning make sure you don't end up taking the log of a negative number).

OpenStudy (anonymous):

thank you :) i got x = +/- 2.5